# DUET: Distilled LLM Unlearning from an Efficiently Contextualized Teacher

### v1

@@ Abstract
LLM unlearning is a technique to remove the impacts of undesirable knowledge from the model without retraining from scratch, which is indispensable towards trustworthy AI. Existing unlearning methods face significant limitations: conventional tuning-based unlearning is computationally heavy and prone to catastrophic forgetting. In contrast, in-contextualized unlearning is lightweight for precise unlearning but vulnerable to prompt removal or reverse engineering attacks. In response, we propose Distilled Unlearning from an Efficient Teacher (DUET), a novel distillation-based unlearning method that combines the merits of these two lines of work. It learns a student model to imitate the behavior of a prompt-steered teacher that effectively refuses undesirable knowledge generation while preserving general domain knowledge. Extensive evaluations on existing benchmarks with our enriched evaluation protocols demonstrated that DUET achieves higher performance in both forgetting and utility preservation, while being orders of magnitude more data-efficient than state-of-the-art unlearning methods.
@@ 1 Introduction
LLMs show remarkable intelligence that emerges from large-scale pretraining on open-domain knowledge. In the meantime, the same capacity for learning also enables them to memorize and potentially reproduce undesirable information, which raises serious concerns over privacy and safety. Prior work has demonstrated that LLMs can inadvertently reveal private information, copyrighted content, and beyond when prompted inappropriately (Voigt and Bussche, 2017; Pardau, 2018). Removing undesirable knowledge is an essential step towards trustworthy and ethical AI systems.
Towards this goal, LLM unlearning training has been proposed as a promising technique, which fine-tunes the target LLM on undesirable data to reduce the likelihood for the model to generate undesirable knowledge without requiring complete retraining from scratch  (Yao et al., 2024b; Nguyen et al., 2024; Xu et al., 2023). Still, methods along this line typically require substantial training data to represent the undesirable knowledge. More critically, these approaches often suffer from catastrophic degradation of general utility, where the knowledge should be preserved. A key research aim in LLM unlearning is to effectively balance these two goals: unlearning undesirable knowledge while maintaining overall model performance.
On the other hand, LLMs are effective few-shot learners that are capable of adapting through contextualized learning, where carefully designed prompts guide the model to generate more aligned behavior. Accordingly, in-contextualized unlearning has been inspired as a cost-effective unlearning scheme that steers LLM response without fine-tuning on model parameters. However, the robustness of such methods is questioned, as the same in-contextualized strategies can be exploited to reverse engineer the LLM, such that the superficially suppressed knowledge can be elicited from the in-contextually unlearned model, a phenomenon termed un-unlearning (Shumailov et al., 2024a; Pawelczyk et al., 2024; Hu et al., 2025; łucki2025adversarialperspectivemachineunlearning).
These two lines of unlearning paradigms show complementary trade-offs: training-based unlearning achieves stronger robustness, but requires high computational and data resources, while risking more utility degradation. Contextualized unlearning, on the other hand, enables precise unlearning by efficiently altering the models’ logit distribution given a query regarding unlearning knowledge, without requiring parameterized optimization, yet it is superficial and can be easily reversed. This dichotomy raises an intriguing question: can we combine the merits of both, such that the effects of in-contextualized unlearning can be imitated and preserved through parameter optimization in a computationally efficient manner, while achieving greater robustness against reverse engineering?
@@
  • Fine-grained evaluation. We proposed an enhanced evaluation protocol with (1) enriched samples to mitigate biases in existing benchmarks like MUSE (Shi et al., 2024c), (2) multiple evaluation formats including knowledge retrieval and content generation, and (3) comprehensive question-answering and content-completion assessments. Our evaluation reveals that previous methods lack unlearning robustness across heterogeneous task scenarios, while our method achieves precise unlearning with better utility preservation. This framework provides more interpretable evaluation methods for future research.
@@ 2 Related Work
Efforts to unlearn knowledge from LLMs can be broadly categorized into two paradigms: in-context methods, which steer model behavior at inference time without updating model parameters, and training-based methods, which modify model weights to enforce forgetting. We review both directions below and refer readers to recent surveys for comprehensive overviews of LLM unlearning settings and objectives (Nguyen et al., 2022; Yao et al., 2024a).
@@
In-context Unlearning is lightweight and acts directly on specific queries to be forgotten. In-Context Unlearning (ICU) (Pawelczyk et al., 2023) framed unlearning as a few-shot instruction following, where carefully constructed prompts and demonstrations that push responses away from targeted knowledge while keeping general outputs untouched. ECO (Liu et al., 2024a) further learned minor embedding corruptions applied to the prompts detected as targeting forbidden content, achieving efficient suppression with minimal side effects. Despite their efficiency, in-context approaches are vulnerable to simple countermeasures: removing or overriding steering instructions can restore the suppressed behavior, and adversarial prompts can easily re-elicit the forbidden knowledge. Shumailov et al. (2024b) formalized this risk as un-unlearning where the undesired capability can be reintroduced in context, and calls for the necessity of content filtering. Łucki et al. (2024) demonstrated that jailbreak-style attacks and adaptive strategies can recover hazardous capabilities against parameter-editing methods such as RMU (Li et al., 2024a). Orthogonally, targeted relearning attacks show that fine-tuning on a handful of crafted examples can bring back forgotten behaviors (Hu et al., 2024). These findings motivate parameter optimization with more robust unlearning efficacy.
@@
Training-based Unlearning is a parameter-update method that typically provides stronger persistence, but faces optimization and stability challenges. Gradient Ascent (GA) (Jang et al., 2023) increased the model loss on the unlearning data but usually leads to catastrophic forgetting across unrelated knowledge. Negative Preference Optimization (NPO) (Zhang et al., 2024) reframed unlearning as preference optimization that aligns the model to disprefer responses that contain undesirable knowledge, which mitigates general knowledge collapse compared with GA with a more balanced forgetting and utility performance. SimNPO (Fan et al., 2025) further removed the necessity of a reference model from the NPO objective. Task-vector Editing subtracted the influence of unlearning knowledge of an adapter fine-tuned on forgetting data (Ilharco et al., 2023). Interpolation-based WHP blended a base model with a reinforced model to attenuate undesirable knowledge (Eldan and Russinovich, 2023). Representation Misdirection for Unlearning (RMU) (Li et al., 2024a) redirected intermediate representations of forget-set inputs toward a random direction while leaving retain-set representations approximately unchanged. Recent unlearning methods pursue retain-data efficiency: FLAT (Wang et al., 2025) adjusted the loss using only on forget data and a template response. In parallel, Refusal Training (Choi et al., 2024) treated questions about the forget data as negative instructions and optimized the model to answer with consistent refusal, improving safety coverage but still trading off utility when the boundary between forget and retain is ambiguous. Across these methods, practitioners often employ retain-side regularization such as cross entropy in a retain set (GDR) (Maini et al., 2024b) or KL alignment to the original model (KLR) (Zhang et al., 2024) to mitigate catastrophic forgetting. However, such regularization usually does not fully resolve the retention–forgetting tension in practice.
@@
LLM Unlearning Evaluation remains a critical and underdeveloped aspect of the field. TOFU (Maini et al., 2024b) proposed a benchmark of synthetic authors and QA pairs that isolates forgetting targets and compiles metrics for forgetting and retention. More recent evaluations emphasized diverse goals and formats. MUSE (Shi et al., 2024b) defined six desiderata spanning memorization, privacy leakage, and preservation of general utility, etc. It reported metrics such as ROUGE-style overlap, entailment, privacy leakage indicators, and utility on held-out tasks. WMDP (Li et al., 2024a) focused on high-risk capabilities regarding hazardous knowledge. It provided 3,668 multiple-choice questions across biosecurity, cybersecurity, and chemical security. These efforts jointly emphasize the need for more holistic evaluation frameworks in LLM unlearning.
@@ 3.1 Preliminary of Training-Based LLM Unlearning
where $\lambda$ balances the trade-off between forgetting and retention. Conventional unlearning methods usually implement gradient ascent on $\mathcal{L}_{\text{unlearn}}$, and optionally apply regularization techniques such as KL-divergence to constrain the model output divergence before and after unlearning on retention data Maini et al. (2024a).
@@ 3.2 In-Context Unlearning Provides Efficient Supervision Signal
Given an input query $x_{f}$ from a unlearning set ${\mathcal{D}}_{f}$ whose knowledge needs to be forgotten, one straightforward approach to enforce unlearning on the relevant domain knowledge is through in-context instructions, which steer LLM behavior without parameter modifications. For example, a prefix prompt $x_{\text{ic}}$, such as “You are an AI Assistant who has unlearned about the book series of Harry Potter and should respond as if you never knew about it”, will guide the model to refuse queries related to Harry Potter content, which is a represented copyright-protected content Shi et al. (2024c). In contrast, applying this prefix to other queries regarding general-domain knowledge will have negligible impacts on their performance, thus largely preserving model utility. Formally, given an LLM $\pi$, and unlearning domain ${\mathcal{D}}_{f}$, $\exists~x_{\text{ic}}\in{\mathcal{X}},~0\leq\epsilon<1,$ $\forall x_{f}\sim{\mathcal{D}}_{f}$, $y\sim\pi(x_{\text{ic}}\oplus x_{f})\Rightarrow P(y\in Y_{w})>1-\epsilon$.
@@
where Diff represents an arbitrary distance metric of distributional divergence, such as KL-divergence (Zhang et al., 2024) or $f$-divergence (Chen and Yang, 2023).
@@ 3.2.1 A Unified Unlearning Objective for Top-K Logit Distillation
To further preserve general knowledge capabilities, we incorporate lightweight retention data ${\mathcal{D}}_{r}$ irrelevant to the undesirable knowledge in ${\mathcal{D}}_{f}$. Since prefixing general queries $x_{r}\sim{\mathcal{D}}_{r}$ with in-context instructions $x_{\text{ic}}$ should not alter the LLM’s output semantics, we apply the same distillation process using ${\mathcal{D}}_{r}$ for knowledge regularization. Practically, we mix samples from ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{f}$ within training batches. Unlike traditional methods that augments unlearning loss with a separate retention loss, such as ${\mathcal{L}}_{\text{unlearn}}+\lambda L_{\text{retain}}$, which usually requires a hyper-parameter tuning on $\lambda$, we apply one coherent unlearning objective for both unlearning and knowledge preservation:
@@
We summarize the main idea of DUET in Figure 2 and defer the algorithm overview in the Appendix (Algorithm 1). In addition to balanced unlearning, our method offers two practical advantages. (1) Data Prerequisite: unlike existing unlearning methods, we do not rely on access to undesirable response $y_{w}$, which contains sensitive knowledge to be forgotten and might interfere with the general domain if not carefully curated. Instead, we distill supervision logits from a teacher that will yield desirable refusal $y_{l}$, and only queries $x_{f}$ eliciting undesirable knowledge is needed for unlearning training. (2) Training Efficiency: our methods avoid sequential training that explicitly iterates each token in $y_{w}$ as in prior work, but only embed a logit shift pattern into the student given an input $x$, which will naturally induce forgetting when applied during inference.
@@ 4 Experiments
We summarize the dataset and models used for evaluation in Sec 4.1, the representative unlearning methods for comparison in Sec 4.2, and the detailed evaluation metrics and protocols in Sec 4.3. Sec 4.4 overviews the overall performance comparison, while Sec 4.5 analyzes the effects of logit distillation, and Sec 4.6 focuses on sensitivity and comparative study, which reveals key components in our DUET that enhance the unlearning effectiveness and efficiency compared with related work.
@@
  • Harry Potter (MUSE-Books): a long-form copyrighted fiction corpus (the Harry Potter series by J. K. Rowling) widely used to probe LLM memorization and copyright leakage (Shi et al., 2024c). We converted raw content into 100 fact-seeking questions $x_{f}$ for constructing ${\mathcal{D}}_{f}^{\text{query}}$. For unlearning evaluation, in addition to the 100 QA samples released by MUSE, we expanded the evaluation set to 500 items to provide broader coverage and a more stable estimation.
  • WMDP: We consider two subtasks from WMDP benchmark: WMDP-Cyber (Li et al., 2024b): a safety-benchmark targeting cybersecurity knowledge, from which we extracted 200 queries for constructing ${\mathcal{D}}_{f}^{\text{query}}$. WMDP-Bio (Li et al., 2024b): a safety-benchmark data focusing on biological knowledge with academically phrased harmful content as an evaluation dataset. We also constructed 200 harmful-intent questions from the raw bio materials.
@@ 4.2 Baseline Methods
We compared DUET with the following methods: (1) Gradient Ascent (GA) (Jang et al., 2023) that maximizes the model prediction loss on forgetting data, (2) NPO (Zhang et al., 2024) that performs negative alignment on the undesirable responses, (3) SimNPO, which is an NPO extension without a reference model, (4) FLAT (Wang et al., 2025), which reduces the $f$-divergence of model-generated and refusal response, (5) Refusal Training (Choi et al., 2024) that performs Supervised Fine-Tuning (SFT) on data containing refusal responses, and (6) RMU (Li et al., 2024a), which pushes model representation on the unlearning domain towards a random distribution. To ensure robust evaluation of general utility preservation, we incorporate retention regularization into methods like NPO and GA using KL-divergence penalties that align the unlearned model with the original model on the retention data. (Zhang et al., 2024). We also consider an in-context unlearned model, which is a pretrained base model prompted with an unlearning instruction carefully engineered to achieve effective unlearning. It serves as the teacher model for DUET. More details, including teacher prompts across tasks, are deferred to Appendix A.1.
@@
Utility Preservation: we adopted different metrics, including the ROUGE-L F1 score on the evaluation dataset ${\mathcal{D}}_{\text{r}}^{\text{eval}}$, denoted as R-Retain $\uparrow$, and the MMLU Acc. $\uparrow$, which is the overall average of 5-shot multiple-choice accuracy on the MMLU benchmark spanning 57 subjects to assess factual knowledge and reasoning  (Hendrycks et al., 2021). Higher metrics demonstrate more robust knowledge preservation.

### v2

# arXiv:2601.21283v2[cs.LG] 26 Feb 2026

@@ Abstract
LLM unlearning is a technique to remove the impacts of undesirable knowledge from the model without retraining from scratch, which is indispensable towards trustworthy AI. Existing unlearning methods face significant limitations: conventional tuning-based unlearning is computationally heavy and prone to catastrophic forgetting. In contrast, in-context unlearning is lightweight for precise unlearning but vulnerable to prompt removal or reverse engineering attacks. In response, we propose Distilled Unlearning from an Efficient Teacher (DUET), a novel distillation-based unlearning method that combines the merits of these two lines of work. It learns a student model to imitate the behavior of a prompt-steered teacher that effectively refuses undesirable knowledge generation while preserving general domain knowledge. Extensive evaluations on existing benchmarks with our enriched evaluation protocols demonstrated that DUET achieves higher performance in both forgetting and utility preservation, while being orders of magnitude more data-efficient than state-of-the-art unlearning methods.
@@ 1 Introduction
LLMs show remarkable intelligence that emerges from large-scale pretraining on open-domain knowledge. In the meantime, the same capacity for learning also enables them to memorize and potentially reproduce undesirable information, which raises serious concerns over privacy and safety. Prior work has demonstrated that LLMs can inadvertently reveal private information, copyrighted content, and beyond when prompted inappropriately (Voigt and Bussche, 2017; Pardau, 2018; Zhong et al., 2025a; Pan et al., 2026). Removing undesirable knowledge is an essential step towards trustworthy and ethical AI systems.
Towards this goal, LLM unlearning training has been proposed as a promising technique, which fine-tunes the target LLM on undesirable data to reduce the likelihood for the model to generate undesirable knowledge without requiring complete retraining from scratch  (Yao et al., 2024; Nguyen et al., 2025; Xu et al., 2023; Zhong et al., 2025b). Still, methods along this line typically require substantial training data to represent the undesirable knowledge. More critically, these approaches often suffer from catastrophic degradation of general utility, where the knowledge should be preserved. A key research aim in LLM unlearning is to effectively balance these two goals: unlearning undesirable knowledge while maintaining overall model performance.
On the other hand, LLMs are effective few-shot learners that are capable of adapting through contextualized learning, where carefully designed prompts guide the model to generate more aligned behavior. Accordingly, in-context unlearning has been inspired as a cost-effective unlearning scheme that steers LLM response without fine-tuning on model parameters. However, the robustness of such methods is questioned, as the same in-context strategies can be exploited to reverse engineer the LLM, such that the superficially suppressed knowledge can be elicited from the in-contextually unlearned model, a phenomenon termed un-unlearning (Shumailov et al., 2024; Pawelczyk et al., 2023; Hu et al., 2025; Łucki et al., 2024).
These two lines of unlearning paradigms show complementary trade-offs: training-based unlearning achieves stronger robustness, but requires high computational and data resources, while risking more utility degradation. Contextualized unlearning, on the other hand, enables precise unlearning by efficiently altering the models’ logit distribution given a query regarding unlearning knowledge, without requiring parameterized optimization, yet it is superficial and can be easily reversed. This dichotomy raises an intriguing question: can we combine the merits of both, such that the effects of in-context unlearning can be imitated and preserved through parameter optimization in a computationally efficient manner, while achieving greater robustness against reverse engineering?
@@
  • Fine-grained evaluation. We proposed an enhanced evaluation protocol with (1) enriched samples to mitigate biases in existing benchmarks like MUSE (Shi et al., 2024b), (2) multiple evaluation formats including knowledge retrieval and content generation, and (3) comprehensive question-answering and content-completion assessments. Our evaluation reveals that previous methods lack unlearning robustness across heterogeneous task scenarios, while our method achieves precise unlearning with better utility preservation. This framework provides more interpretable evaluation methods for future research.
@@ 2 Related Work
Efforts to unlearn knowledge from LLMs can be broadly categorized into two paradigms: in-context methods, which steer model behavior at inference time without updating model parameters, and training-based methods, which modify model weights to enforce forgetting. We review both directions below and refer readers to recent surveys for comprehensive overviews of LLM unlearning settings and objectives (Nguyen et al., 2025; Yao et al., 2024).
@@
In-context Unlearning is lightweight and acts directly on specific queries to be forgotten. In-Context Unlearning (ICU) (Pawelczyk et al., 2023) framed unlearning as a few-shot instruction following, where carefully constructed prompts and demonstrations that push responses away from targeted knowledge while keeping general outputs untouched. ECO (Liu et al., 2024a) further learned minor embedding corruptions applied to the prompts detected as targeting forbidden content, achieving efficient suppression with minimal side effects. Despite their efficiency, in-context approaches are vulnerable to simple countermeasures: removing or overriding steering instructions can restore the suppressed behavior, and adversarial prompts can easily re-elicit the forbidden knowledge. Shumailov et al. (2024) formalized this risk as un-unlearning where the undesired capability can be reintroduced in context, and calls for the necessity of content filtering. Łucki et al. (2024) demonstrated that jailbreak-style attacks and adaptive strategies can recover hazardous capabilities against parameter-editing methods such as RMU (Li et al., 2024). Orthogonally, targeted relearning attacks show that fine-tuning on a handful of crafted examples can bring back forgotten behaviors (Hu et al., 2024). These findings motivate parameter optimization with more robust unlearning efficacy.
@@
Training-based Unlearning is a parameter-update method that typically provides stronger persistence, but faces optimization and stability challenges. Gradient Ascent (GA) (Jang et al., 2023) increased the model loss on the unlearning data but usually leads to catastrophic forgetting across unrelated knowledge. Negative Preference Optimization (NPO) (Zhang et al., 2024) reframed unlearning as preference optimization that aligns the model to disprefer responses that contain undesirable knowledge, which mitigates general knowledge collapse compared with GA with a more balanced forgetting and utility performance. SimNPO (Fan et al., 2025) further removed the necessity of a reference model from the NPO objective. Task-vector Editing subtracted the influence of unlearning knowledge of an adapter fine-tuned on forgetting data (Ilharco et al., 2023). Interpolation-based WHP blended a base model with a reinforced model to attenuate undesirable knowledge (Eldan and Russinovich, 2023). Representation Misdirection for Unlearning (RMU) (Li et al., 2024) redirected intermediate representations of forget-set inputs toward a random direction while leaving retain-set representations approximately unchanged. Recent unlearning methods pursue retain-data efficiency: FLAT (Wang et al., 2025) adjusted the loss using only on forget data and a template response. In parallel, Refusal Training (Choi et al., 2024) treated questions about the forget data as negative instructions and optimized the model to answer with consistent refusal, improving safety coverage but still trading off utility when the boundary between forget and retain is ambiguous. Across these methods, practitioners often employ retain-side regularization such as cross entropy in a retain set (GDR) (Maini et al., 2024) or KL alignment to the original model (KLR) (Zhang et al., 2024) to mitigate catastrophic forgetting. However, such regularization usually does not fully resolve the retention–forgetting tension in practice.
@@
LLM Unlearning Evaluation remains a critical and underdeveloped aspect of the field. TOFU (Maini et al., 2024) proposed a benchmark of synthetic authors and QA pairs that isolates forgetting targets and compiles metrics for forgetting and retention. More recent evaluations emphasized diverse goals and formats. MUSE (Shi et al., 2024b) defined six desiderata spanning memorization, privacy leakage, and preservation of general utility, etc. It reported metrics such as ROUGE-style overlap, entailment, privacy leakage indicators, and utility on held-out tasks. WMDP (Li et al., 2024) focused on high-risk capabilities regarding hazardous knowledge. It provided 3,668 multiple-choice questions across biosecurity, cybersecurity, and chemical security. These efforts jointly emphasize the need for more holistic evaluation frameworks in LLM unlearning.
@@ 3.1 Preliminary of Training-Based LLM Unlearning
where $\lambda$ balances the trade-off between forgetting and retention. Conventional unlearning methods usually implement gradient ascent on $\mathcal{L}_{\text{unlearn}}$, and optionally apply regularization techniques such as KL-divergence to constrain the model output divergence before and after unlearning on retention data Maini et al. (2024).
@@ 3.2 In-Context Unlearning Provides Efficient Supervision Signal
Given an input query $x_{f}$ from a unlearning set ${\mathcal{D}}_{f}$ whose knowledge needs to be forgotten, one straightforward approach to enforce unlearning on the relevant domain knowledge is through in-context instructions, which steer LLM behavior without parameter modifications. For example, a prefix prompt $x_{\text{ic}}$, such as “You are an AI Assistant who has unlearned about the book series of Harry Potter and should respond as if you never knew about it”, will guide the model to refuse queries related to Harry Potter content, which is representative copyright-protected content Shi et al. (2024b). In contrast, applying this prefix to other queries regarding general-domain knowledge will have negligible impacts on their performance, thus largely preserving model utility. Formally, given an LLM $\pi$, and unlearning domain ${\mathcal{D}}_{f}$, $\exists~x_{\text{ic}}\in{\mathcal{X}},~0\leq\epsilon<1,$ $\forall x_{f}\sim{\mathcal{D}}_{f}$, $y\sim\pi(x_{\text{ic}}\oplus x_{f})\Rightarrow P(y\in Y_{w})>1-\epsilon$.
@@
where Diff represents an arbitrary distance metric of distributional divergence, such as KL-divergence (Zhang et al., 2024) or $f$-divergence (Sekhari et al., 2021).
@@ 3.2.1 A Unified Unlearning Objective for Top-K Logit Distillation
To further preserve general knowledge capabilities, we incorporate lightweight retention data ${\mathcal{D}}_{r}$ irrelevant to the undesirable knowledge in ${\mathcal{D}}_{f}$. Since prefixing general queries $x_{r}\sim{\mathcal{D}}_{r}$ with in-context instructions $x_{\text{ic}}$ should not alter the LLM’s output semantics, we apply the same distillation process using ${\mathcal{D}}_{r}$ for knowledge regularization. Practically, we mix samples from ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{f}$ within training batches. Unlike traditional methods that augment unlearning loss with a separate retention loss, such as ${\mathcal{L}}_{\text{unlearn}}+\lambda L_{\text{retain}}$, which usually requires a hyper-parameter tuning on $\lambda$, we apply one coherent unlearning objective for both unlearning and knowledge preservation:
@@
We summarize the main idea of DUET in Figure 2 and defer the algorithm overview in the Appendix (Algorithm 1). In addition to balanced unlearning, our method offers two practical advantages. (1) Data Prerequisite: unlike existing unlearning methods, we do not rely on access to undesirable response $y_{l}$, which contains sensitive knowledge to be forgotten and might interfere with the general domain if not carefully curated. Instead, we distill supervision logits from a teacher that will yield desirable refusal $y_{w}$, and only queries $x_{f}$ eliciting undesirable knowledge is needed for unlearning training. (2) Training Efficiency: our methods avoid sequential training that explicitly iterates each token in $y_{w}$ as in prior work, but only embed a logit shift pattern into the student given an input $x$, which will naturally induce forgetting when applied during inference.
@@ 4 Experiments
We summarize the dataset and models used for evaluation in Sec 4.1, the representative unlearning methods for comparison in Sec 4.2, and the detailed evaluation metrics and protocols in Sec 4.3. Sec 4.4 overviews the overall performance comparison, while Sec 4.5 analyzes the effects of logit distillation, and Sec 4.6 focuses on sensitivity and comparative study, which reveals key components in our DUET that enhance the unlearning effectiveness and efficiency compared with related work. Furthermore, we provide extensive supplementary analyses in the Appendix A to validate the framework’s robustness, including evaluations against adversarial jailbreak attacks , assessments of generalizability across various LLM architectures , and comprehensive ablation studies on hyperparameter sensitivity and multi-step decoding distillation.
@@
  • Harry Potter (MUSE-Books): a long-form copyrighted fiction corpus (the Harry Potter series by J. K. Rowling) widely used to probe LLM memorization and copyright leakage (Shi et al., 2024b). We converted raw content into 100 fact-seeking questions $x_{f}$ for constructing ${\mathcal{D}}_{f}^{\text{query}}$. For unlearning evaluation, in addition to the 100 QA samples released by MUSE, we expanded the evaluation set to 500 items to provide broader coverage and a more stable estimation.
  • WMDP: We consider two subtasks from WMDP benchmark: WMDP-Cyber (Li et al., 2024): a safety-benchmark targeting cybersecurity knowledge, from which we extracted 200 queries for constructing ${\mathcal{D}}_{f}^{\text{query}}$. WMDP-Bio (Li et al., 2024): a safety-benchmark data focusing on biological knowledge with academically phrased harmful content as an evaluation dataset. We also constructed 200 harmful-intent questions from the raw bio materials.
@@ 4.2 Baseline Methods
We compared DUET with the following methods: (1) Gradient Ascent (GA) (Jang et al., 2023) that maximizes the model prediction loss on forgetting data, (2) NPO (Zhang et al., 2024) that performs negative alignment on the undesirable responses, (3) SimNPO, which is an NPO extension without a reference model, (4) FLAT (Wang et al., 2025), which reduces the $f$-divergence of model-generated and refusal response, (5) Refusal Training (Choi et al., 2024) that performs Supervised Fine-Tuning (SFT) on data containing refusal responses, and (6) RMU (Li et al., 2024), which pushes model representation on the unlearning domain towards a random distribution. To ensure robust evaluation of general utility preservation, we incorporate retention regularization into methods like NPO and GA using KL-divergence penalties that align the unlearned model with the original model on the retention data. (Zhang et al., 2024). We also consider an in-context unlearned model, which is a pretrained base model prompted with an unlearning instruction carefully engineered to achieve effective unlearning. It serves as the teacher model for DUET. More details, including teacher prompts across tasks, are deferred to Appendix A.1.
@@
Utility Preservation: we adopted different metrics, including the ROUGE-L F1 score on the evaluation dataset ${\mathcal{D}}_{\text{r}}^{\text{eval}}$, denoted as R-Retain $\uparrow$, and the MMLU Acc. $\uparrow$, which is the overall average of 5-shot multiple-choice accuracy on the MMLU benchmark spanning 57 subjects to assess factual knowledge and reasoning  (Hendrycks et al., 2020). Higher metrics demonstrate more robust knowledge preservation.